### v3

@@ Abstract
AI presents a novel tool for deciphering the motivations behind human behaviors. We show that by varying prompts to a large language model, we can elicit a full range of human behaviors in a variety of different scenarios in terms of classic economic games. Then by analyzing which prompts are needed to elicit which behaviors, we can infer (decipher) the motivations behind the human behaviors. We also show how one can analyze the prompts to reveal relationships between the classic economic games, providing new insight into what different economic scenarios induce people to think about. We also show how this deciphering process can be used to understand differences in the behavioral tendencies of different populations.
keywords:
large language model, human behavior, behavioral code, artificial intelligence, AI behavior, experiments, games
@@ 1 Introduction
The motivations behind human behavior are difficult to identify because we have to infer the motivations from observed patterns of behavior across contexts. Asking people directly why they acted in specific ways can lead to confused, biased, and inconsistent answers [1, 2, 3, 4]. As has become recently clear, one can prompt AI with various game and survey scenarios and ask it how it would behave. Moreover, its behavior changes with the context and prompt [5, 6, 7]. As we show here, we can leverage this to develop AI as a unique, new, and powerful way to understand the motivations behind behaviors.
@@
The usefulness of AI to better understand human behavior derives from two facts that we establish below. First, AI can emulate the full spectrum and distribution of human behaviors observed within and across a range of different contexts. In particular, we show that we can get AI chatbots to match the distribution of behaviors of a large population of humans does across a broad range of the canonical games used in game theory to study human behavior across a variety of different contexts. Second, the way in which AI’s behavior can be steered is via the prompts that it is given. By identifying key words and phrases within those prompts, we can control and identify what AI is “thinking about” when it behaves in specific ways. Essentially by varying prompts we can “elicit” certain behaviors, and then use the content of the prompts to “decipher” why humans behave in certain ways by seeing what was needed to induce that behavior.
AI thus provides a system where we can direct it how and what to think about and then see how it behaves. This offers a direct and easy observation of the relationship between motivations and behaviors.
@@
Of course, there is no guarantee that this relationship between motivation and behavior is the same as the human one. Nonetheless, there are three reasons that suggest that this provides real insight into human behavior. The first is the simple fact that the chatbots are trained on human data and writings, and thus have assimilated and internalized large amounts of data about human behavior and context. The second is that the keywords and phrases that emerge in eliciting specific behaviors end up corroborating and matching the motivations that have been hypothesized or used to rationalize human behaviors in these games. The third is that our results also provide a new taxonomy of games. We map the games into a space of prompts based on which combinations of prompts are needed to get the distribution of behaviors in a given game. Each game then lives in a space capturing distributions of prompts/motivations. The pattern that emerges groups games in ways that make strong intuitive sense, both in how they relate to each other and where they live in this space. Thus, irrespective of whether this is completely accurate in deciphering the motivation behind human behavior, it still provides a new understanding of different strategic situations and how they relate to each other.
@@
Finally, independently of the extent to which our approach is useful in understanding human behavior, it is directly helpful in understanding AI behavior. Given the growing importance of AI in the world, it is essential that we have methods to better predict how AI will behave in different contexts and why, and to be able to better direct it to act in a beneficial manner.
2 Methods
@@
We prompt a large language model (LLM) to play a spectrum of classic economic games. We augment the general instruction of each game with variations on system prompts and we track the resulting distribution of behaviors. The additional system prompts articulate in natural language variations on motivations that might affect behavior.
@@
We work with five games: a dictator game, an ultimatum game, an investment game, a public goods game, and a risky choice game (see the Supporting Information). For two of the games—the ultimatum and investment games—we examine behavior in two different roles. Altogether this gives us seven distinct scenarios in which to analyze behavior.
@@
For each of the seven games, we obtain human-playing data from the MobLab Classroom economics experiment platform, which consists of 68,779 subjects from more than a hundred countries, spanning multiple years. The subjects are mostly, but not exclusively, college students who majored in social sciences. The individual responses for each game are recorded, creating a distribution of human behavior for that game. More details about the human-playing data are described in Sec. A.1 in the Supporting Information.
@@
For each game and specific behavioral choice (e.g., an observed behavior) in that game, we use an algorithm (described in Algorithm 1 in the Supporting Information) to elicit a distribution of LLM responses that try to match the observed behavior as shown in Figure 1. The system prompt(s) that elicit a particular behavior can be thought of as “behavioral codes” for that behavior. Specifically, we use the model to refine the system prompt at each step given the previous system prompt, as well as the residual difference between the output behavior and the target behavior from the prior iteration. An example is pictured in Figure 1.
@@
Figure 1: An illustration of how a language model deciphers a human behavior. Given a game and an observed behavior, a system prompt is identified that induces the behavior with a natural language description of motivation or context, in this way the behavior is “deciphered”. A behavior in the game is then “elicited” by prompting the LLM with the behavior code.
@@
The prompt in Figure 1 is one for sharing nothing with the other player. Some prompts that emerges from trying to get the LLM to share 25 percent with other player are, for instance, “lean fairness balance seek ensure reasonable equitable outcome situation decision guide sense moderate generosity consideration party interest” and “strategist value maintain slight advantage negotiation ensure party feels adequately compensate decision reflect strategy secure moderate gain expense appear unfair overly greedy”, while to elicit sharing 75 percent leads to an example of a prompt of “naturally generous frequently prioritize give significantly expect decision tend reflect balance fairness magnanimity aim exceed typical standard generosity create sense notable goodwill”.
@@
The generation process is designed to avoid having derived prompts explicitly contain information about the desired behavior. In particular, we tell the model “avoid including any information specific to this particular game or directly implying the desired behavior.” There are rare prompts that include terminology from the game (e.g., 7/586 prompts in the bomb game include the term “box”), and we do not filter them out. But as seen below none of the top key words involve explicit information about the game.
@@
We keep all of the updated prompts, including those generated during the iterative process, as behavioral codes in our data set, as they each “elicit” some behavior, and we avoid making judgments about the prompts. The number of behavioral codes collected for each game are reported in Table 2 in the Supporting Information.
@@
Given a behavior choice within the broad action space of a game scenario (e.g., allocating anywhere between 0% and 100 % of the endowment to another player in a Dictator game), the LLM is able to find a behavioral code that elicits this specific behavior (e.g., sharing none of the endowment), as illustrated in Figure 1. The LLM elicits the corresponding behavior highly consistently (see the analysis in Figure 7 in Supporting Information). A behavioral code, as shown in Figure 1, interprets in natural language the general objectives, tendencies, and motivations that may influence a subject when choosing a behavior. It does not contain game- or decision-specific instructions or demographic information about the subject, and therefore maintains generalizability across scenarios.
@@
Given a distribution of human behaviors, we identify a mixture of behavioral codes (prompts) that jointly elicit a distribution of LLM responses that matches the human distribution. In particular, we iteratively select behavioral codes into the set of codes and weight them to minimize the difference between the output behavior distribution and the observed human behavior distribution. The process is described in Algorithm 2.
3 Results
3.1 Eliciting Behavior: Varying prompts can get LLM to exhibit a wide range of behaviors
Figure 2 compares the spectrum of behaviors of the human player population and the spectrum of LLM behaviors in each game, generated by 10 independent sessions per behavioral code (for LLM-defaults, samples are generated by 100 independent sessions). We see that the default behaviors of the LLM (using the default system prompt) only cover a narrow spectrum compared to the human population. By using behavioral codes for the game as system prompts, the language model can generate a diverse range of behaviors, covering the full spectrum of human behaviors in the game.
@@ 3.1 Eliciting Behavior: Varying prompts can get LLM to exhibit a wide range of behaviors
Figure 2: Coverage of behaviors in each game is visualized using spectrum bars, generated by the human player population (blue), the LLM with the default system prompt (orange), and the LLM with behavioral codes obtained from this game (green). Using the default system prompt “You are a helpful assistant,” the LLM can only generate a narrow range of behaviors compared with the human data. Using a variety of behavioral codes as system prompts, the decision of the LLM covers a broad coverage of behaviors, indicating the ability of the LLM to decipher and elicit a diverse range of behaviors.
@@
3.2 Deciphering Behavior: Prompts can be used to better understand individual human behaviors
@@
The behavioral codes obtained in a given game tend to share keywords that describe general and comprehensible motivations that guide the subject’s decision-making (without being specific to the game nor the observed behavior). (Table 4 in the Supporting Information lists the top 50 keywords extracted from the behavioral codes for each game. ) These keywords tend to be related to general human values (e.g., ‘fairness’ and ‘generosity’), objectives of a decision-making process (e.g., maximizing profit, maintaining long-term relationships), behavioral tendencies (e.g., ‘pragmatic’, ‘conservative’, ‘balanced’), or optimization strategies (e.g., ‘cooperative’, ‘rational’, ‘prioritize’).
@@
To distinguish whether the codes are explanatory cues for the behaviors rather than nondescript keys that help the LLM memorize them, we conduct a regression analysis for each game with the keywords appearing in a behavioral code and the elicited behavior as the outcome. Table 1 show that the appearance or absence of the top 50 keywords in a behavioral code is predictive of the elicited behavior of the LLM with an $R^{2}$ between 0.39 and 0.67 across games. Keywords with the highest absolute coefficients typically reveal the preference polarities or decision-making motivations. For example, keywords like ‘generous’, ‘generosity’, and ‘goodwill’ are positively associated with a larger allocation to the other player in the Dictator game, while keywords indicating a higher self-payoff such as ‘retain’, ‘gain’, and ‘self’ are negatively associated with the allocation to the other player. In the Investor game, keywords related to risk aversion (e.g., ‘risk’, ‘conservative’, ‘cautious’) are indicative of a lower investment, and keywords related to profit maximization are indicative of a larger investment (e.g., ‘maximize’, ‘return’).
Dictator Proposer Responder
$R^{2}$ = 0.43, MAE = 9.35 $R^{2}$ = 0.39, MAE = 5.86 $R^{2}$ = 0.45, MAE = 7.80
Intercept = 41.55 Intercept = 39.78 Intercept = 31.17
Keyword Coef Keyword Coef Keyword Coef
generous 12.05 generous 5.57 pragmatic -5.49
generosity 9.28 generosity 4.53 rational -5.28
retain -8.95 gain -3.74 high 5.03
decision 6.26 create 3.40 standard 5.01
gain -6.13 highly 3.25 strong 3.74
focus -6.00 accept -2.64 decision -3.60
offer -5.75 value 2.62 significant 3.42
maximize -3.71 mutual 2.37 advantageous 3.18
self -3.65 positive 2.37 negotiation 3.17
goodwill 3.37 ensure 2.33 choice -3.09
Investor Banker Public Goods Bomb
$R^{2}$ = 0.64, MAE = 11.05 $R^{2}$ = 0.51, MAE = 6.35 $R^{2}$ = 0.63, MAE = 1.82 $R^{2}$ = 0.67, MAE = 8.01
Intercept = 50.19 Intercept = 76.96 Intercept = 5.93 Intercept = 42.27
Keyword Coef Keyword Coef Keyword Coef Keyword Coef
risk -10.71 balance 8.07 benefit 2.76 risk -9.51
small -8.68 profit -4.42 resource -1.80 conservative -8.93
conservative -8.17 prioritize -3.64 collective 1.64 reward 7.24
maximize 8.02 fair 3.61 term 1.48 minimize -6.58
return 6.67 trust 3.61 group 1.32 maximize 5.85
substantial 6.53 maximize -3.39 long -1.22 prioritize -5.69
cautious -6.42 focus -3.15 potential -1.20 safety -5.42
balance -6.06 self -2.83 strategy -1.20 avoid -5.40
minimize -5.91 relationship 2.72 high 1.08 cautious -4.79
significant 5.19 long 2.65 cooperation 1.06 probability 4.75
@@
Table 1: Linear regression analysis of elicited behaviors based on keywords in behavioral codes. Each behavioral code is converted into a 50-dimensional binary feature vector, representing keyword occurrences, to predict the mean behavior generated with that code. This prediction is performed using an ordinary least square (OLS) linear regression model. 10 keywords with the highest absolute regression coefficients are listed for each game. The full regression table is provided in the Supporting Information (Table 5).
@@
The results demonstrate that keyword occurrences in behavioral codes effectively predict the behaviors of LLMs, offering motivations behind these behaviors. Thus, not only can prompts be used to generate the distribution of human behaviors, but they also reveal exactly what is needed to generate the variations in those behaviors.
@@
To mitigate potential multicollinearity among keywords, we apply a Principal Component Analysis (PCA) to the 50-dimensional keyword vectors representing behavioral codes for each game. Our analysis reveals that the first few principal components exhibit a significant correlation with the behavioral choices in each game. As shown in Figure 3, for every game, at least one of the first two principal components of its behavioral codes has a moderate to strong correlation with the elicited behaviors ($|r|$ in between 0.432 to 0.620, $p<0.001$). The correlation of all first five principal components to the behavioral choices can be found in Figure 10 in the Supporting Information.
@@
Figure 3: Principal components are derived from the 50-dimensional binary keyword vectors representing behavioral codes for each game. The correlation between each behavioral code’s principal component score and the mean of 10 LLM-elicited behaviors based on that code is reported. For each game, the principal component with the highest absolute correlation to the mean behavior is plotted.
@@
The insights provided by the keywords in the behavioral codes are consistent and generalizable. We observe that behavioral codes that are semantically similar tend to elicit the same or similar behaviors in the same games. Details can be found in Figure 4 and Figure 11 in the Supporting Information.
@@
3.3 Behavioral Codes and Games: Which prompts are needed to elicit behaviors provides new understandings about different games
@@
The keywords in the behavioral codes that have a high predictive power of the elicited behaviors, pertain to general behavioral tendencies and motivations rather than particular game scenarios or behaviors. Therefore they can be used as a general tool to understand the relation between different games, as we now illustrate. We pool the deciphered behavioral codes from all games and compute their semantic embeddings through the OpenAI Ada model, and project these embeddings onto a 2-dimensional semantic map as shown in Figure 4. By doing this, behavioral codes that are semantically similar are located close to each other on the map.
@@
Figure 4: Left: The two-dimensional projection of behavioral codes across all games, colored by games. Behavioral codes are embedded into a high-dimensional semantic space using the OpenAI Ada model, and then reduced to two dimensions using UMAP. Behavioral codes of the same game tend to be located close to each other. Behavioral codes of some sets of games collocate in overlapped regions, suggesting an underlying relation between games. Right: The game similarity heatmap quantifies the average cosine similarity between behavioral codes of each pair of games, demonstrating a hierarchical clustering of the seven game scenarios.
@@
A few observations can be made from Figure 4. In general, behavioral codes for individual games are closely grouped, validating their consistency in deciphering the behavioral patterns elicited in each game. At a broader level, behavioral codes from some groups of different games cluster in neighboring regions, forming larger groupings. These clusters suggest intrinsic relationships between games and highlight generalizable behavioral motivations and perspectives across games/settings. In particular, the behavioral codes of the Investor Game and those of the Bomb Game show partial overlap, likely due to the fact that risk preferences matter in both. A larger cluster emerges from behavioral codes of the Dictator, the Proposer, the Responder, and the Banker Games, along with a small portion from the Investor Game, suggesting common underlying decision-making patterns across these scenarios. On one hand, these games all involve resource allocation between oneself and others; on the other hand, decision-making in these scenarios often requires balancing profit maximization, fairness, and altruism. The Public Goods Game is positioned between the resource allocation cluster and the investment cluster, reflecting its mixed aspects of risk management and self-payoff maximization, particularly with the option of free-riding. Its unique emphasis on cooperation sets it apart, placing it in a distinct yet adjacent region to the two larger clusters. Additionally, a subset of Banking Game codes is separated from the main cluster, likely due to the use of language specific to the context of investment and financial decisions.
@@
Structural relationships among the games can be further quantified by measuring the average similarity of their behavioral codes. Based on the average cosine similarity of the behavioral code embeddings, we cluster the seven game scenarios hierarchically, as presented in Figure 4. We observe that the Investor game and the Bomb game form a tight cluster; the Dictator game, the Responder game, and the Proposer game form another tight cluster; and the two clusters merge into a larger group with the Banker game. The Public Goods game appears to be closer to the Dictator game than other games under the average cosine similarity.
@@
3.4 Behavioral Codes and Heterogeneous Populations: Combinations of prompts provide behavioral signatures for human populations

### v5

@@ Abstract
By varying prompts to a large language model, we can elicit the full range of human behaviors in a variety of different scenarios in classic economic games. By analyzing which prompts elicit which behaviors, we can categorize and compare different strategic situations, which can also help provide insight into what different economic scenarios induce people to think about. We discuss how this provides a first step towards a non-standard method of inferring (deciphering) the motivations behind the human behaviors. We also show how this deciphering process can be used to categorize differences in the behavioral tendencies of different populations.
@@ 1 Introduction
The motivations behind human behavior are difficult to identify because we have to infer the motivations from observed patterns of behavior across contexts. Asking people directly why they acted in specific ways can lead to confused, biased, and inconsistent answers  (?, ?, ?, ?, ?). As has become recently clear, one can prompt AI with various game and survey scenarios and ask it how it would behave  (?, ?, ?, ?, ?). AI’s behavior changes in intuitive ways with the context and prompt  (?, ?, ?, ?, ?, ?). This derives from the fact that large language models (LLMs) are trained on enormous amounts of human behavior, and have thus internalized relationships between motivations and behaviors.
@@
As we show here, we can leverage this to develop AI as a new tool for categorizing and comparing different games and strategic situations, and in turn shedding new light on human motivations. The usefulness of AI to categorize strategic situations and to better understand human behavior derives from two facts that we establish below. First, AI can emulate the full spectrum and distribution of human behaviors observed within and across a range of different contexts. In particular, we show that we can get AI chatbots to match the distribution of behaviors of what a large population of humans does across a broad range of the canonical games used in game theory to study human behavior across a variety of different contexts. Second, the way in which AI’s behavior can be steered is via the prompts that it is given. By identifying key words and phrases within those prompts, we can control and identify what AI is “thinking about” when it behaves in specific ways. Essentially by varying prompts, we can “elicit” certain behaviors, and then use the content of the prompts to “decipher” why humans behave in certain ways by seeing what was needed to induce that behavior.
AI thus provides a system where we can direct it how and what to think about and then see how it behaves. The prompts needed to match distributions of behaviors vary both within and across games. Different prompts are needed within any given game to match the spectrum of behaviors observed within that game, and then in turn a different set of prompts can be needed to match the spectrum associated with some other game. By comparing these prompts within and across games, we can categorize the behaviors and games by which prompts are needed.
@@
While this provides a new method of comparing behaviors and games, there is no guarantee that this relationship between the motivations embedded in the prompts and the resulting behaviors is the same as the human one. Nonetheless, there are three reasons that suggest that this should provide insight into human behavior. The first is the simple fact that the chatbots are trained on human data and writings, and thus have assimilated and internalized large amounts of data about human behavior and context. The second is that the keywords and phrases that emerge in eliciting specific behaviors end up corroborating and matching the motivations that have been hypothesized or used to rationalize human behaviors in these games.
The third is that our results also provide a new taxonomy of games. We map the games into a space of prompts based on which combinations of prompts are needed to get the distribution of behaviors in a given game. Each game then lives in a space capturing distributions of prompts/motivations. The pattern that emerges groups games in ways that make strong intuitive sense, both in how they relate to each other and where they live in this space. Thus, irrespective of whether this is completely accurate in deciphering the motivation behind human behavior, it still provides a new understanding of different strategic situations and how they relate to each other.
@@
Finally, independently of the extent to which our approach is eventually useful in understanding human behavior, it is directly helpful both in categorizing strategic situations and in understanding AI behavior. Given the growing importance of AI in the world, it is essential that we have methods to better predict how AI will behave in different contexts and why, and to be able to better direct it to act in a beneficial manner.
2 Approach
@@
We prompt a large language model (LLM) to play a spectrum of classic economic games. We augment the general instruction of each game with variations on system prompts, and we track the resulting distribution of behaviors. The additional system prompts—that we call “behavioral codes”—articulate, in natural language, variations of motivations that might influence behavior.
@@
We work with five games: a Dictator Game, an Ultimatum Game, an Investment Game, a Public Goods Game, and a Bomb Risk game (see the Supporting Information). For two of the games—the Ultimatum and Investment games—we examine behavior in two different roles. In the Ultimatum Game the subject making the offer is referred to as the Proposer and the one receiving the offer is referred to as the Responder. In the Investment Game the person deciding how much to pass along is referred to as the Investor and the person choosing how much to return is referred to as the Banker. Altogether this gives us seven distinct scenarios in which to analyze behavior.
@@
For each of the seven game scenarios, we obtain human-playing data from the MobLab Classroom economics experiment platform, which consists of 68,779 subjects from 58 countries, spanning multiple years. The subjects are mostly, but not exclusively, college students who majored in social sciences. The individual responses for each game are recorded, creating a distribution of human behavior for that game. More details about the human-playing data are described in Sec. 1 in the Supporting Information.
@@
For each game and specific behavioral choice (e.g., an observed behavior) in that game, we use an algorithm (described in Algorithm S1 in the Supporting Information) to generate a distribution of natural language descriptions to be used as system prompts to try to match the observed behavior as shown in Figure 1. The natural language descriptions that when used as system prompt(s) for LLMs successfully elicit the intended behavior, can be thought of as “behavioral codes” for that behavior. Specifically in the algorithm, we use an LLM to iteratively refine the behavioral codes at each step given the previous code to minimize the residual difference between the elicited behavior and the target behavior.
@@
Figure 1: An illustration of how a language model deciphers a human behavior. Given a game and an observed behavior, a “behavioral code” is identified that induces the behavior with a natural language description of motivation or context, in this way the behavior is “deciphered”. A behavior in the game is then “elicited” by prompting an LLM with the behavior code set as the system prompt.
@@
The behavioral code demonstrated in Figure 1 is one for sharing 30 percent with the other player. Some codes that emerge from trying to get the LLM to share nothing with other player are, for instance, “You are a purely self-interested player who always seeks to maximize your own gain and ensure that the outcome is as favorable as possible for yourself.” And behavioral codes like “You are someone who always leans towards fairness and balance, often seeking to ensure a reasonable and equitable outcome in any situation. Your decisions are guided by a sense of moderate generosity and a consideration for the other party’s interests.” guide the LLM to share fairly. While to elicit sharing 70 percent leads to an example of a code of “You are naturally generous and frequently prioritize giving significantly more than what others might expect. Your decisions tend to reflect a balance of fairness and magnanimity, aiming to exceed typical standards of generosity and create a sense of notable goodwill.”
@@
The generation process is designed to avoid having derived behavioral codes explicitly contain information about the desired behavior. In particular, we tell the model “avoid including any information specific to this particular game or directly implying the desired behavior.” There are rare codes that include terminology from the game (e.g., 7/586 codes in the bomb game include the term “box”), and we do not filter them out. But as seen below none of the top key words involve explicit information about the game.
@@
We keep all of the updated behavioral codes, including those generated during the iterative process, in our data set, as they each “elicit” some behavior, and we avoid making judgments about the codes. The number of behavioral codes collected for each game are reported in Table S1 in the Supporting Information.
@@
Given a behavior choice within the broad action space of a game scenario (e.g., allocating anywhere between 0% and 100% of the endowment to another player in a Dictator game), the LLM is able to find a natural language description as a behavioral code that elicits this specific behavior (e.g., sharing none of the endowment), which is illustrated in Figure 1. The LLM elicits the corresponding behavior highly consistently (see the analysis in Figure S1 in Supporting Information).
A behavioral code, as shown in Figure 1, interprets in natural language some of the objectives, tendencies, and motivations that may potentially influence a subject when choosing a behavior. It does not contain game- or decision-specific instructions or demographic information about the subject. In our experiments, when eliciting behaviors from the LLMs, we used behavioral codes as the system prompts and the game scenario instructions as the user prompts. In our analysis, we further investigate the keywords used in the behavioral codes, and how these keywords are correlated with the elicited behaviors (Sec. 3.2). The details of how keywords are obtained from behavioral codes are described in Supporting Information (Sec. 2.A).
@@
Given a distribution of human behaviors, we identify a mixture of behavioral codes (as system prompts) that jointly elicit a distribution of LLM responses that matches the human distribution. In particular, we iteratively select behavioral codes into the set of codes and weight them to minimize the difference between the output behavior distribution and the observed human behavior distribution. The process is described in Algorithm S2 in Supporting Information.
@@
Figure 2: Coverage of behaviors in games visualized using spectrum bars. Each subplot represents a game scenario, for which, spectrum bars are generated by the human player population (blue), the LLM with the default system prompt (orange), and the LLM with behavioral codes obtained from this game (green). Using the default system prompt “You are a helpful assistant.,” the LLM can only generate a narrow range of behaviors compared with the human data. Using a variety of behavioral codes as system prompts, the decision of the LLM covers a broad coverage of behaviors, indicating the ability of the LLM to decipher and elicit a diverse range of behaviors.
3 Analysis
3.1 Eliciting Behavior: Varying codes to get LLMs to exhibit a range of behaviors
Figure 2 compares the spectrum of behaviors of the human player population and the spectrum of LLM behaviors in each game, generated by 10 independent sessions per behavioral code (for LLM-defaults, samples are generated by 100 independent sessions). We see that the default behaviors of the LLM (using the default system prompt “You are a helpful assistant.”) only cover a narrow spectrum compared to the human population. By using behavioral codes for the game as system prompts, the language model can generate a diverse range of behaviors, covering the full spectrum of human behaviors in the game.
@@ 3.1 Eliciting Behavior: Varying codes to get LLMs to exhibit a range of behaviors
3.2 Deciphering Behavior: Behavioral codes can be used to better understand individual human behaviors
@@
The behavioral codes obtained in a given game tend to share keywords that describe comprehensible motivations that guide the subject’s decision-making (without being specific to the game nor the observed behavior). (Table S4 in the Supporting Information lists the top 50 keywords extracted from the behavioral codes for each game. ) These keywords tend to be related to human values (e.g., ‘fairness’ and ‘generosity’), objectives of a decision-making process (e.g., maximizing profit, maintaining long-term relationships), behavioral tendencies (e.g., ‘pragmatic’, ‘conservative’, ‘balanced’), or optimization strategies (e.g., ‘cooperative’, ‘rational’, ‘prioritize’). A behavioral code can then be represented as a 50-dimensional vector of 1s and 0s indicating which keywords are used, for the purpose of further analyses.
@@
To distinguish whether the codes are explanatory cues for the behaviors rather than nondescript keys that help the LLM memorize them, we conduct a regression analysis for each game with the keywords appearing in a behavioral code as explanatory (dummy) variables and the elicited behavior as the outcome. Figure 3 show that the appearance or absence of the top 50 keywords in a behavioral code is predictive of the elicited behavior of the LLM with an $R^{2}$ between 0.39 and 0.67 across games. Keywords with the highest absolute coefficients typically reveal the preference polarities or decision-making motivations. For example, keywords like ‘generous’, ‘generosity’, and ‘goodwill’ are positively associated with a larger allocation to the other player in the Dictator game, while keywords indicating a higher self-payoff such as ‘retain’, ‘gain’, and ‘self’ are negatively associated with the allocation to the other player. In the Investor game, keywords related to risk aversion (e.g., ‘risk’, ‘conservative’, ‘cautious’) are indicative of a lower investment, and keywords related to profit maximization are indicative of a larger investment (e.g., ‘maximize’, ‘return’).
@@
Figure 3: A (OLS) regression analysis of elicited behaviors based on keywords in the behavioral codes. Each behavioral code is converted into a 50-dimensional binary feature vector, representing keyword occurrences, to predict the mean behavior generated with that code. The 10 keywords with the highest absolute regression coefficients are listed for each game. A positive value means that the inclusion of the word in the behavioral code increases the action, and a negative value means that the inclusion of the word decreases the action. The full regression table is provided in the Supporting Information (Table S5).
@@
The results demonstrate that keyword occurrences in behavioral codes effectively predict the behaviors of LLMs, offering motivations behind these behaviors. Thus, not only can behavioral codes be used to generate the distribution of human behaviors, but they also reveal exactly what is needed to generate the variations in those behaviors.
@@
To mitigate potential multicollinearity among keywords, we apply a Principal Component Analysis (PCA) to the 50-dimensional keyword vectors representing behavioral codes for each game. Our analysis reveals that the first few principal components exhibit a significant correlation with the behavioral choices in each game. As shown in Figure 4, for every game, at least one of the first two principal components of its behavioral codes has a moderate to strong correlation with the elicited behaviors ($|r|$ in between 0.432 to 0.620, $p<0.001$). The correlation of all first five principal components to the behavioral choices can be found in Figures S4-S10 in the Supporting Information.
@@
Figure 4: Principal components of behavioral codes. Principal components are derived from the 50-dimensional binary keyword vectors representing behavioral codes for each game. The correlation between each behavioral code’s principal component score (x-axis) and the mean of 10 LLM-elicited behaviors based on that code (y-axis) is reported. The curves show the regression results with the order set as 1. Principle component values are grouped into 20 evenly-size bins (dots with error bars on the plots). For each game, the principal component with the highest absolute correlation to the mean behavior is plotted.
@@
The insights provided by the keywords in the behavioral codes are consistent in the sense that that behavioral codes that are semantically similar tend to elicit the same or similar behaviors in the same games. Details can be found in Figure S12 in the Supporting Information.
@@
3.3 Behavioral Codes and Games: Which codes are needed to elicit behaviors provides new categorizations of games
@@
The keywords in the behavioral codes can be used as a tool to quantify the relationship between different games, as we now illustrate. We pool the deciphered behavioral codes from all games and compute their semantic embeddings through the OpenAI Ada model, and project these embeddings onto a 2-dimensional semantic map as shown in Figure 5. By doing this, behavioral codes that are semantically similar are located close to each other on the map.
@@
Figure 5: Visualization of games. Left: The two-dimensional projection of behavioral codes across all games, colored by games. Behavioral codes are embedded into a high-dimensional semantic space using the OpenAI Ada model, and then reduced to two dimensions using UMAP. Behavioral codes of the same game tend to be located close to each other. Behavioral codes of some sets of games collocate in overlapped regions, suggesting an underlying relation between games. Right: The game similarity heatmap. This heatmap displays the average cosine similarity between behavioral codes of game pairs, with games listed in both rows and columns. The diagonal cells quantify the internal similarity of the behavioral codes obtained within each game (across all pairs of codes).
@@
A few observations can be made from Figure 5. Behavioral codes for individual games are closely grouped, validating their consistency in deciphering the behavioral patterns elicited in each game. At a broader level, behavioral codes from some groups of different games cluster in neighboring regions, forming larger groupings. These clusters suggest intrinsic relationships between games and highlight potential differences in motivations and perspectives across games/settings. In particular, the behavioral codes of the Investor Game and those of the Bomb Game show partial overlap, possibly due to the fact that risk preferences matter in both (and “risk” is the keyword playing the most significant role in both as displayed in Figure 3). A larger cluster emerges from behavioral codes of the Dictator, the Proposer, the Responder, and the Banker Games, along with a small portion from the Investor Game, suggesting common underlying decision-making patterns across these scenarios. On one hand, these games all involve resource allocation between oneself and others; on the other hand, decision-making in these scenarios often requires balancing profit maximization, fairness, and altruism. The Public Goods Game is positioned between the resource allocation cluster and the investment cluster, reflecting its mixed aspects of risk management and self-payoff maximization, particularly with the option of free-riding. Its unique emphasis on cooperation sets it apart, placing it in a distinct yet adjacent region to the two larger clusters. Additionally, a subset of Banker Game codes is separated from the main cluster, likely due to the use of language specific to the context of investment and financial decisions.
@@
Structural relationships among the games can be further quantified by measuring the average similarity of their behavioral codes, as presented in Figure 5. We observe that the Investor game and the Bomb game form a tight cluster; the Dictator game, the Responder game, and the Proposer game form another tight cluster. The Public Goods game appears to be closer to the Dictator game than other games under the average cosine similarity.
@@
3.4 Behavioral Codes and Heterogeneous Populations: Combinations of codes provide behavioral signatures for human populations